# Writer-Defined AI Personas for On-Demand Feedback Generation

### v1

@@ 5.3.4. Impact of the Generated Feedback
Participants found that the feedback helped them and had an impacted on their text. We also observed impacts on the writers themselves.
@@
Most participants (12 people) made changes to their texts in out editor based on feedback they received. For instance, $P_{6}$ added a personal anecdote to their text after reading that the feedback suggested this. At the end, they said “I ended up making a lot of adjustments based off the feedback.” Related, 9 out of 11 participants in Study 2 reported in the Likert questions that the feedback influenced their writing (see Section 5.1.2).
@@
Another finding was that we observed participants anthropomorphizing the system through their personas. For instance, $P_{16}$ made a sad face after being told that the writing time was almost up, saying “But I didn’t talk to ”best friend” yet” before requesting feedback from that persona. Participants referenced their personas as if they could be real individuals. For example, they said “Let’s see what X thinks about this”, where X is the name of their persona ($P_{6}$, $P_{12}$). However, some participants (5 people) gave their personas generic names like “Persona 1”.
@@ 5.4. How well can contemporary LLMs support AI personas? (RQ3)
Overall, we found that it is possible to use a contemporary LLM (GPT-3.5) to support writers with AI personas, albeit with some limitations and room for improvement, related to model and/or prompt engineering.
@@ 5.4.2. Quality of the Feedback
The feedback was longer than needed. For example, $P_{8}$ said “it took me a while to read like the actual feedback”. $P_{10}$ and $P_{15}$ suggested a “bullet point” format in contrast to the fully spelled out comments that were generated. Related to this verbosity, the feedback was partly perceived as unspecific, such as “very generic”’ ($P_{9}$) or “too abstract” ($P_{10}$). For example, $P_{10}$ said that “there was truth in there, but it was like a bit drowned out by the things that surrounded it”. And $P_{11}$ said that “It just tells me I have to do this and it doesn’t come up with the specific examples”. $P_{12}$ wished for a “follow-up feature where you can […] ask it again, for example, just something specific”.
@@ 6.1. Taxonomy of AI Personas for Feedback
Based on our study, we provide a taxonomy of personas and which dimension users considered. The most frequently observed attribute that users took into account was expertise. In the academic setting we investigated, this was frequently related to professional roles and qualifications. Such expertise was also defined by mentioning specific, well-known institutions like the Oxford University ($P_{10}$) or the New York Times ($P_{13}$). $P_{15}$ used the abstract term “scientific writer”, while $P_{11}$ referred to a top researcher by name.
Moreover, social relation was an important consideration. This is reflected in personas named “best friend”, “my colleague”, and “my manager”. One participant even mentioned that it might be fun to define a persona such as “mom” ($P_{7}$). We also observed that the valence of personas was adapted. Some users considered a “worst-case” person like a “mean reviewer” ($P_{12}$), others emulated their “PhD advisor” ($P_{14}$) or their “PhD colleague” ($P_{11}$), as more positively supportive personas. Another dimension we identified was the level of involvement, again ranging from closely related personas, such as “my manager” ($P_{7}$) and “my PhD colleague” ($P_{11}$), to less involved personas such as a “New York Times journalist” ($P_{13}$).
@@
Our study sample was mostly young academics. We expect that other populations will define different concrete personas. While the presented taxonomy is on a higher level of abstraction, insights from other user groups might extend or refine it. Future work could explore this further with different samples.
@@ 6.2.2. Writers use AI personas to consider multiple perspectives
One key benefit that we identified is that AI personas in our prototype empowered users to compare multiple perspectives to reflect on their writing. $P_{6}$ reported several occasions where adding different perspectives made him “step outside of [their] writing and look at it”, which helped them notice things they had not noticed before. $P_{9}$ compared the AI persona approach to the video game Sims, where the user can “create a lot of characters”. Strategically, writers added further personas throughout their writing time in the study (Figure 5). Overall, participants created a diverse range of personas (Section 5.3.1, Section A.4).
@@ 6.2.3. Writers use AI personas to gain access to perspectives on demand
The AI personas provided users with on-demand feedback from the perspectives of people that they otherwise would not have been able to receive feedback from. Participants liked the convenience of getting “instant feedback” ($P_{14}$). $P_{13}$ also appreciated getting “feedback from these different kinds of people without actually having to find them”. $P_{7}$ described how they could create an avatar of professors that they have not met yet, based on a description of their backgrounds and research interests. While we rarely observed such advanced use of AI personas to “simulate” people a user has not yet met, it shows the potential of the concept and provides ample opportunities for future work.
@@ 6.3.1. Writers find it difficult to define personas
Initially defining a persona was challenging because it was difficult for participants to articulate the key aspects of the target group. Nevertheless, they found several strategies, such as thinking of a specific (real) person, using our provided example attributes (Figure 2), and iteratively refining a persona after reading its feedback. Future research should explore further ways of supporting the creation of AI personas. For example, one idea is to provide a library of personas that allow writers to collaboratively refine and reuse personas. This is akin to prompt collections for LLM-based systems, where users share the instructions that they empirically found to be particularly helpful for a given task. We discuss another direction in Section 6.5.3.
@@ 6.3.2. Generated feedback can be verbose, repetitive and unspecific
The other key issue that we identified concerns text qualities. Participants found the feedback too verbose and – connected to this – repetitive and unspecific. One reason for this was that feedback started by repeating persona information. We discuss further prompt engineering as a solution strategy in Section 6.5.2. Conceptually, based on participants’ comments and iterative refinement of personas, we expect tradeoffs between feedback length and helpfulness as well as the transparency of the persona’s “argument” (feedback is this because of that persona information). Future work could use our insights as a starting point to explore the “sweet spot” here. This might also be dynamic. For example, as some comments related to repetitiveness suggest, it might be useful to reduce feedback length over time as users become familiar with a persona.
@@ 6.4. The Representativity of Contemporary AI Systems
As AI personas here combine user-defined personas and LLMs, they might reproduce stereotypes. Participants did not comment on feedback being stereotypical. However, they criticized that it unnecessarily repeated attributes. While seeing how the LLM links personal attributes and feedback in its written output might be insightful (at least initially), it is also a likely place where stereotypes could surface.
Our findings should be considered in light of the dangers of LLMs described by Bender et al. (2021). They provide a broad overview of stereotypical and derogatory associations that LLMs infer from their training data, including sensitive attributes like gender, race, ethnicity, and disability status. Bender et al. argue that reliance on internet data is associated with a risk of encoding the dominant view and failing to represent changing social views after training (“value-lock”). They also explain why the mere size of the training corpus does not guarantee diversity since only a certain privileged group can fully participate online. This also connects to work that discussed which human values are represented by LLMs and which are not (Johnson et al., 2022).
@@
In view of these aspects, we consider it essential that the concept of AI personas, especially when put into practice, is pursued with careful consideration of how the resulting representations of people can be facilitated and evaluated. For example, systems that provide AI personas could warn users of the potentially limited representativity of contemporary AI systems. One key part of our concept is that writers can change the personas at any time. While that is not the same as having control over the underlying LLM, it provides one pathway of influence to the user, which might be helpful in cases where writers themselves recognize generated output as unsuitable.
Looking ahead, we anticipate that these questions will not remain limited to the design of specific tools, but likely impact HCI research more broadly. For example, it seems increasingly likely to receive AI-generated responses when conducting survey-based user research online (Hämäläinen et al., 2023).
@@ 6.5.1. Shifting the design focus: From human-AI to writer-reader relationships
This socio-technical expansion of the design perspective on AI writing tools complements recent work on the social dynamics of writing support (Gero et al., 2023), as well as calls for such tools to consider social aspects of language (Kulkarni and Raheja, 2023). It also contributes balance to design perspectives derived from the writing model by Flower and Hayes (1981), which is increasingly considered in the design of writing tools (see (Bhat et al., 2023; Gero et al., 2022)) but focuses on processes within the writer and thus has little to say by itself about the perspectives of other affected people, such as readers. Finally, we expect this framing to be fruitful beyond writing, for example, to also keep in mind the creator-audience relationship when designing tools for creators of images, videos, podcasts or games.
@@ 6.5.2. Prompt engineering
The presented findings are based on one model and one way of prompting it. Thus, we do not claim to present the “optimal” implementation of AI personas for feedback. While further design iterations were beyond the scope of this paper, we discuss possible future directions.
Future work could explore how prompting techniques  (e.g. (Brown et al., 2020; Wei et al., 2022)) can be leveraged to make the LLM output less verbose. Kabir et al. (2023) found LLMs to provide verbose answers to software engineering queries, underlining that this might be an issue of broader interest for interactive use of such models. For our use case, further prompts might also be explored with regard to how representative the feedback is (see Section 6.4).
Underlying these considerations is the question emerging in HCI of how to support end users in prompting effectively (Zamfirescu-Pereira et al., 2023). Here, the persona concept could be explored as a framing that allows non-experts to steer a model towards a desired role through an understandable UI and metaphor. In a way, this makes explicit that there are (customizable) UI concepts to design around the “system instruction” prompts of systems like ChatGPT (cf. Section 2.2). This direction also connects to recent work on configurable UIs for LLMs (Kim et al., 2023b, a).
More broadly, with prompt engineering and LLMs, building functional prototypes of AI tools is becoming increasingly feasible within user-centred design. In light of such a process, we see our work as a starting point towards understanding AI personas in various use cases. We hope others will join us to make them even more helpful for writers and other creators.
@@ 6.5.3. Connecting writers and readers through reader-defined personas
Given the complexity of defining a persona, it is essential to think of ways of supporting this further. One avenue could be reader-defined personas. Rather than asking individual writers to imagine personas that represent their readers, future work could ask readers to define their personas themselves. This would not only simplify the workflow for writers. Following the mantra “Nothing About Us Without Us”, that scholars of Critical Disability Studies in HCI subscribe to (Spiel et al., 2020), it would allow readers to articulate what they need and explicate what the AI-based persona should pay attention to. Conceptually, the AI would then serve both stakeholders (readers and writers) as a mediating executable proxy perspective. Future work could explore this as an important kind of computer-supported cooperative work that socio-technical interventions can facilitate.
@@ 7. Conclusion
While AI feedback comes with limitations (e.g. lack of specificity, cf. (Gero et al., 2023)), it offers enticing benefits, such as direct editor integration, almost instant turnarounds, and infinite availability and patience with a struggling writers’ repeated requests. We look forward to future work towards the larger vision of leveraging AI to make us write better, to consider diverse reader perspectives, and to achieve a writing goal that is inherently human: sharing ideas, through language, that others can understand.
@@ A.3. Likert Results
Shows the Likert data as horizontal diverging bar charts. Overall, participants appreciated the provided features. Most found that the feedback was helpful and had an influence on their work. They still perceived themselves as the authors. Participants were critical of the feedback length (too long). Finally, our sample had a high interest in AI.

### v2

@@ 5.3.4. Impact of the Generated Feedback
Participants found that the feedback helped them and had an impact on their text. We also observed impacts on the writers themselves.
@@
Most participants (12 people) made changes to their texts in our editor based on the feedback they received. For instance, $P_{6}$ added a personal anecdote to their text after reading that the feedback suggested this. At the end, they said “I ended up making a lot of adjustments based off the feedback.” Related, 9 out of 11 participants in Study 2 reported in the Likert questions that the feedback influenced their writing (see Section 5.1.2).
@@
Another finding was that we observed participants anthropomorphizing the system through their personas. For instance, $P_{16}$ made a sad face after being told that the writing time was almost up, saying “But I didn’t talk to ‘best friend’ yet” before requesting feedback from that persona. Participants referenced their personas as if they could be real individuals. For example, they said “Let’s see what X thinks about this”, where X is the name of their persona ($P_{6}$, $P_{12}$). However, some participants (5 people) gave their personas generic names like “Persona 1”.
@@ 5.4. How well can contemporary LLMs support AI personas? (RQ3)
Overall, we found that it is possible to use a contemporary LLM (GPT-3.5) to support writers with AI personas, albeit with some limitations and room for improvement, related to model and prompt engineering.
@@ 5.4.2. Quality of the Feedback
The feedback was longer than needed. For example, $P_{8}$ said “it took me a while to read like the actual feedback”. $P_{10}$ and $P_{15}$ suggested a “bullet point” format in contrast to the fully spelled out comments that were generated. Related to this verbosity, the feedback was partly perceived as unspecific, such as “very generic” ($P_{9}$) or “too abstract” ($P_{10}$). For example, $P_{10}$ said that “there was truth in there, but it was like a bit drowned out by the things that surrounded it”. And $P_{11}$ said that “It just tells me I have to do this and it doesn’t come up with the specific examples”. $P_{12}$ wished for a “follow-up feature where you can […] ask it again, for example, just something specific”.
@@ 6.1. Taxonomy of AI Personas for Feedback
Based on our study, we provide a taxonomy of personas and which dimension users considered. The most frequently observed attribute that users took into account was expertise. In the specific academic setting we investigated, this was frequently related to competencies, professional roles and qualifications. Such expertise was also defined by mentioning specific, well-known institutions like Oxford University ($P_{10}$) or the New York Times ($P_{13}$). $P_{15}$ used the abstract term “scientific writer”, while $P_{11}$ referred to a top researcher by name.
Moreover, social relation was an important consideration. This is reflected in personas named “best friend”, “my colleague”, and “my manager”. One participant even mentioned that it might be fun to define a persona such as “mom” ($P_{7}$). We also observed that the valence of personas was taken into account. Some users considered a “worst-case” person like a “mean reviewer” ($P_{12}$), others emulated their “PhD advisor” ($P_{14}$) or their “PhD colleague” ($P_{11}$), as more positively supportive personas. Another dimension we identified was the level of involvement, again ranging from closely related personas, such as “my manager” ($P_{7}$) and “my PhD colleague” ($P_{11}$), to less involved personas such as a “New York Times journalist” ($P_{13}$).
@@
Our study sample was mostly young academics. We expect that other populations will define different concrete personas. While the presented taxonomy is on a higher level of abstraction, insights from other user groups might extend or refine it. Future work could explore this further with different samples. The taxonomy may also have to be adapted for other writing styles. For instance, expertise may play a less important role in the context of fiction writing. Future work could explore how the different factors must be weighted based on genre.
@@ 6.2.2. Writers use AI personas to consider multiple perspectives
One key benefit that we identified is that AI personas in our prototype empowered users to compare multiple perspectives to reflect on their writing (cf. goal 4 in Section 3.1.2). $P_{6}$ reported several occasions where adding different perspectives made him “step outside of [their] writing and look at it”, which helped them notice things they had not noticed before. $P_{9}$ compared the AI persona approach to the video game Sims, where the user can “create a lot of characters”. Strategically, writers added further personas throughout their writing time in the study (Figure 6). Overall, participants created a diverse range of personas (Section 5.3.1, Section A.4).
@@ 6.2.3. Writers use AI personas to gain access to perspectives on demand
The AI personas provided users with on-demand feedback from the perspectives of people that they otherwise would not have been able to receive feedback from (cf. goal 1 in Section 3.1.2). Participants liked the convenience of getting “instant feedback” ($P_{14}$). $P_{13}$ also appreciated getting “feedback from these different kinds of people without actually having to find them”. $P_{7}$ described how they could create an avatar of professors that they have not met yet, based on a description of their backgrounds and research interests. While we rarely observed such advanced use of AI personas to “simulate” people a user has not yet met, it shows the potential of the concept and provides ample opportunities for future work.
@@ 6.3.1. Writers find it difficult to define personas
Initially defining a persona was challenging because it was difficult for participants to articulate the key aspects of the target group, even with the improved UI in study 2 (see Section 3.1.1 and goal 2 in Section 3.1.2). Nevertheless, they found several strategies, such as thinking of a specific (real) person, using our provided example attributes (Figure 2), and iteratively refining a persona after reading its feedback (cf. goal 3 in Section 3.1.2). Future research should explore further ways of supporting the creation of AI personas. For example, one idea is to provide a library of personas that allow writers to collaboratively refine and reuse personas. This is akin to prompt collections for LLM-based systems, where users share the instructions that they empirically found to be particularly helpful for a given task. We discuss another direction in Section 6.5.3.
@@ 6.3.2. Generated feedback can be verbose, repetitive and unspecific
The other key issue that we identified concerns text qualities. Participants found the feedback too verbose and – connected to this – repetitive and unspecific, although our UI enabled them to request feedback for specific text passages, not the whole text (cf. goal 5 in Section 3.1.2). One reason for this was that feedback started by repeating persona information. We discuss further prompt engineering as a solution strategy in Section 6.5.2. Conceptually, based on participants’ comments and iterative refinement of personas, we expect tradeoffs between feedback length and helpfulness as well as the transparency of the persona’s “argument” (feedback is this because of that persona information). Future work could use our insights as a starting point to explore the “sweet spot” here. This might also be dynamic. For example, as some comments related to repetitiveness suggest, it might be useful to reduce feedback length over time as users become familiar with a persona.
@@ 6.4. The Representativity of Contemporary AI Systems
As AI personas are a combination of user-defined personas and LLMs, they have the potential to reproduce harmful stereotypes. In our in-depth qualitative analysis, we did not identify any instances where participants perceived the feedback as perpetuating stereotypes. However, we identified situations where they criticized that the system unnecessarily repeated certain attributes. While seeing how the LLM links personal attributes and feedback in its written output might be insightful (at least initially), it is also a likely place where stereotypes could surface.
Even though we did not encounter stereotypes in practice, the potential dangers of LLMs in this regard have been described in prior work, for instance, by Bender et al. (2021). Prior work provides a broad overview of stereotypical and derogatory associations that LLMs infer from their training data, including sensitive attributes like gender, race, ethnicity, and disability status. Bender et al. argue that reliance on internet data is associated with a risk of encoding the dominant view and failing to represent changing social views after training (“value-lock”). They also explain why the mere size of the training corpus does not guarantee diversity since only a certain privileged group can fully participate online. This connects to work that discussed which human values are represented by LLMs and which ones are not (Johnson et al., 2022).
@@
In view of these aspects, we consider it essential that the concept of AI personas, especially when put into practice, is pursued with careful consideration of how the resulting representations of people can be facilitated and evaluated. Our analysis of the generated feedback did not reveal derogatory or stereotypical statements, such as sexist or racist ones (cf. Section 5.1.3). Nevertheless, systems that provide AI personas could warn users of the potentially limited representativity of contemporary AI systems and how they can arise from limitations in the training data and the training processes.
One key part of our concept is that writers decide on what attributes to define for their personas, which they can change anytime. While that is not the same as having control over the underlying LLM, it provides one pathway of influence to the user, which might be helpful in cases where writers themselves recognize generated output as unsuitable.
That said, there might also be cases where the model’s bias is more subtly expressed in the feedback, such that writers may not immediately notice this. Suppose stereotypes like a sexist statement are detected in generated comments in a research setting like our study. In that case, this may be addressed by sending an educational debrief to participants, as done in prior work (Jakesch et al., 2023) (e.g., informing participants about the incorrect or biased statement, with a link/content from a credible source to counter it). In production systems, regular audits may be necessary to ensure that systems are as free of harmful stereotypes as possible (Sandvig et al., 2014). For this, it may be helpful to inform users what stereotypes were detected, how they were detected, and how they were addressed.
Looking ahead, we anticipate that these questions will not remain limited to the design of specific tools but likely impact HCI research more broadly. For example, it seems increasingly likely to receive AI-generated responses when conducting survey-based user research online (Hämäläinen et al., 2023).
@@ 6.5.1. Shifting the design focus: From human-AI to writer-reader relationships
Related work on explainable AI (XAI) revealed interesting perspectives by recognizing XAI as “socially-situated” (Ehsan et al., 2021). In our context, our socio-technical expansion of the design perspective on AI writing tools complements recent work on the social dynamics of writing support (Gero et al., 2023), as well as calls for such tools to consider social aspects of language (Kulkarni and Raheja, 2023). It also contributes balance to design perspectives derived from the writing model by Flower and Hayes (1981), which is increasingly considered in the design of writing tools (see (Bhat et al., 2023; Gero et al., 2022)) but focuses on processes within the writer and thus has little to say by itself about the perspectives of other affected people, such as readers. Finally, we expect this framing to be fruitful beyond writing, for example, to also keep in mind the creator-audience relationship when designing tools for creators of images, videos, podcasts or games.
@@ 6.5.2. Prompt engineering
The presented findings are based on one model and one way of prompting it. Thus, we do not claim to present the “optimal” implementation of AI personas for feedback. While further design iterations were beyond the scope of this paper, we have started to explore two strategies to address verbosity, following the insights of our study: One strategy is to modify the prompt template to emphasize more strongly that feedback should be concise. In addition, we shortened the text in the few-shot examples. Anecdotally, first observations indicate that this may not be enough to keep generated comments concise consistently. As another strategy, we experimented with applying the model a second time to its own output, using a second prompt that instructs it to make the given feedback more concise. This seemed promising in our first observations and is in line with “chaining” in other prompt-based systems (cf. (Wu et al., 2022)).
Future work could explore in detail how such prompting techniques  (e.g. (Brown et al., 2020; Wei et al., 2022)) can be leveraged specifically to make LLM output less verbose, since this is a broader issue beyond our use case: For example, Kabir et al. (2023) found LLMs to provide verbose answers to software engineering queries, and Sun et al. (2023b) diagnosed that LLMs struggle with hard constraints, including on length (“numerical planning”). For our use case, further prompts might also be explored with regard to how representative the feedback is (see Section 6.4).
Underlying these considerations is the emerging question in HCI of how to support end users in prompting effectively (Zamfirescu-Pereira et al., 2023). Here, the persona concept could be explored as a framing that allows non-experts to steer a model towards a desired role through an understandable UI and metaphor. In a way, this makes explicit that there are (customizable) UI concepts to design around the “system instruction” prompts of systems like ChatGPT (cf. Section 2.2). This direction also connects to recent work on configurable UIs for LLMs (Kim et al., 2023b, a).
More broadly, with prompt engineering and LLMs, building functional prototypes of AI tools is becoming increasingly feasible in user-centred design. In this light, we see our work as a starting point towards understanding AI personas in various use cases. We hope others will join us to make them even more helpful for writers and other creators.
@@ 6.5.3. Connecting writers and readers through reader-defined personas
Given the complexity of defining a persona, it is essential to think of ways of supporting this further. One direction could be reader-defined personas. Rather than asking individual writers to imagine personas that represent their readers, future work could ask readers to define their personas themselves. This would not only simplify the workflow for writers. Following the mantra “Nothing About Us Without Us”, that scholars of Critical Disability Studies in HCI subscribe to (Spiel et al., 2020), it would allow readers to articulate what they need and explicate what the AI-based persona should pay attention to. Conceptually, the AI would then serve both stakeholders (readers and writers) as a mediating “executable proxy” for communication: For example, the writer could refer to the AI persona in moments where direct communication with the readers is not possible, or use it to get feedback more frequently than what is otherwise feasible. Future work could explore this as an important kind of computer-supported cooperative work that socio-technical interventions can facilitate.
@@ 7. Conclusion
While AI feedback comes with limitations (e.g. lack of specificity, cf. (Gero et al., 2023)), it offers enticing benefits, such as direct editor integration, almost instant turnarounds, and infinite availability and patience with a struggling writer’s repeated requests. We look forward to future work towards the larger vision of leveraging AI to make us write better, to consider diverse reader perspectives, and to achieve a writing goal that is inherently human: sharing ideas, through language, that others can understand.
To facilitate such future work, we release our prototype and material here:
@@ A.1. Prompt Template for Generating Feedback
The attribute-description pairs defined by the writers were inserted as indicated by the Python statements above. For example, if a participant defined a persona with the pairs “role: reviewer”, “writing style: formal”, “sentence length: short”, and “occupation: CS professor”, and selected a text “Lorem ipsum dolor sit amet”, then the corresponding part of the prompt template above would concretely look like this:
    Input:
    Text: "Lorem ipsum dolor sit amet"
    Persona:
    - Role: {"role": "reviewer"}
    - Background: {"occupation": "CS professor"}
    - Style: {"writing style": "formal", "sentence length": "short"}
    - Content: {}
@@ A.3. Likert Results
Shows the Likert data as horizontal diverging bar charts. Overall, participants appreciated the provided features. Most found that the feedback was helpful and had an influence on their work. They still perceived themselves as the authors. Participants were critical of the feedback length (too long). Finally, our sample had a high interest in Artificial Intelligence.